# GLASS Flows: Transition Sampling for Alignment of Flow and Diffusion Models

### v1

@@ 1 Introduction
Flow matching and diffusion models have revolutionized the generation of images, videos, and many other data types (Lipman et al., 2022; Albergo et al., 2023; Liu et al., 2022; Song et al., 2020b; Ho et al., 2020). They convert Gaussian noise into realistic images or videos by simulating an ordinary or stochastic differential equation (ODE/SDE). Trained on large web-scale datasets, these models can generate highly realistic images or videos at unprecedented quality. Due to diminishing returns of pre-training these models, many recent works propose methods to improve models at inference-time, i.e. by optimizing additional objectives commonly referred to as rewards (Uehara et al., 2025). These reward alignment algorithms allow to enhance text-to-image alignment (Zhang et al., 2025), solve inverse problems (Chung et al., 2022; He et al., 2023), and improve molecular design (Li et al., 2025). However, as these algorithms achieve higher performance at the expense of more compute, efficiency remains a major challenge for deploying reward alignment algorithms.
@@
$p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t})=\mathbb{P}[X_{t^{\prime}}=x_{t}|X_{t}=x_{t}],\quad(x_{t},x_{t^{\prime}}\in\mathbb{R}^{d},0\leq t<t^{\prime}\leq 1)$ (1)
where $p_{t^{\prime}|t}$ is called the transition kernel. Many reward alignment algorithms rely on sampling from $p_{t^{\prime}|t}$. For example, search algorithms use samples $X_{t^{\prime}}\sim p_{t^{\prime}|t}(\cdot|x_{t})$ as branches of a search tree (for ODE sampling, there would be only one branch). This creates a dilemma: So far, it is not known how to obtain samples $X_{t^{\prime}}\sim p_{t^{\prime}|t}(\cdot|x_{t})$ using ODEs. Therefore, one has to switch from ODE to SDE sampling, losing efficiency in order to use an algorithm that is meant to increase it.
@@
In this work, we present a method to sample from transitions $p_{t^{\prime}|t}$ using ODEs: Gaussian Latent Sufficient Statistic (GLASS) Flows. GLASS Flows combine (1) the high efficiency of ODEs with (2) the controllable stochastic evolution characteristic of SDEs. GLASS Flows construct an “inner flow matching model” to sample from $p_{t^{\prime}|t}$ (see fig. 1). Crucially, this inner flow matching can be easily obtained from pre-trained flow matching models without any fine-tuning using the concept of a sufficient statistic, a fundamental tool in theoretical statistics (Fisher, 1922). Hence, GLASS Flows are a simple plug-in for any algorithm relying on SDE sampling. We apply GLASS Flows to reward alignment and improve the state-of-the-art performance in text-to-image generation. To summarize, we make the following contributions:
@@
    Plugging GLASS Flows into Sequential Monte Carlo and reward guidance procedures for text-to-image generation, we show significant performance improvements at zero cost.
@@ 2 Background and Motivation
In this section, we introduce the necessary background on flow and diffusion models. We follow the flow matching framework (Lipman et al., 2022; Albergo et al., 2023; Liu et al., 2022), yet everything applies similarly to diffusion models. We denote data points with vectors $z\in\mathbb{R}^{d}$ and the data distribution with $p_{\rm{data}}$. Here, $t=0$ corresponds to noise ($\mathcal{N}(0,I_{d})$) and $t=1$ corresponds to data $(p_{\rm{data}})$. To noise data $z\in\mathbb{R}^{d}$, we use a Gaussian conditional probability path $p_{t}(x_{t}|z)$:
@@
where $\alpha_{t},\sigma_{t}\geq 0$ are schedulers with $\alpha_{0}=\sigma_{1}=0$ and $\alpha_{1}=\sigma_{0}=1$ and $\alpha_{t}$ (resp. $\sigma_{t}$) strictly monotonically increasing (resp. decreasing). Making $z\sim p_{\rm{data}}$ random, this induces a corresponding marginal probability path $p_{t}(x_{t})=\mathbb{E}_{z\sim p_{\rm{data}}}[p_{t}(x_{t}|z)]$ which interpolates Gaussian noise $p_{0}=\mathcal{N}(0,I_{d})$ and data $p_{1}=p_{\rm{data}}$. Flow matching models learn the marginal vector field $u_{t}(x_{t})$:
@@ 3 Motivation: Efficient Transitions for Reward Alignment
Sequential Monte Carlo (SMC) methods (Singhal et al., 2025; Skreta et al., 2025) use a transition kernel $p_{t^{\prime}|t}$ as a proposal distribution. Specifically, they evolve $K$ particles $x_{t}^{k}$ via
@@
Search methods (Li et al., 2025; Zhang et al., 2025) consider DDPM sampling as a rollout of a search tree with branches coming from samples from $p_{t^{\prime}|t}$. Beyond sampling branches of the search tree, search methods use approximations of the value function (Li et al., 2025) defined via
@@
As sampling from $p_{1|t}(z|x_{t})$ is only possible with the SDE so far and therefore inefficient, most search methods use approximations of this function (Li et al., 2025; Zhang et al., 2025). Similarly, approximations of the value function $V_{t}(x_{t})$ can also be used to define potentials in SMC procedures.
@@ 4 GLASS Flows
we get samples from the transition kernel $p_{t^{\prime}|t}$ at $s=1$ for an appropriate initial distribution $p_{\text{init}}$. We present a simple algorithm to obtain $u_{s}(\bar{x}_{s}|x_{t},t)$ that we explain in this section (see algorithm 1).
@@
2:  $u\leftarrow{\color[rgb]{0,0,0}u_{t}(x)}$ $\triangleright$ neural net call
3:  return $\frac{1}{\dot{\alpha}_{t}\,\sigma_{t}-\alpha_{t}\,\dot{\sigma}_{t}}(\sigma_{t}u-\dot{\sigma}_{t})$
@@
6:  $S(\mathbf{x})\leftarrow\frac{\mu^{T}\Sigma^{-1}}{\mu^{T}\Sigma^{-1}\mu}[x_{t},\bar{x}_{s}]^{T}$
7:  $t^{*}\leftarrow g^{-1}((\mu^{T}\Sigma^{-1}\mu)^{-1})$
8:  return $D(\alpha_{t^{\star}}S(\mathbf{x}),t^{\star})$
9:
@@
11:  $\mu\leftarrow[\,\alpha_{t},\,\bar{\alpha}_{s}+\bar{\gamma}\,\alpha_{t}\,]^{\top}$
12:  $\Sigma\leftarrow\begin{bmatrix}\sigma_{t}^{2}&\sigma_{t}^{2}\bar{\gamma}\\ \sigma_{t}^{2}\bar{\gamma}&\bar{\sigma}_{s}^{2}+\bar{\gamma}^{2}\sigma_{t}^{2}\end{bmatrix}$
@@
14:  $w_{1}\leftarrow(\partial_{s}\bar{\sigma}_{s})/\bar{\sigma}_{s}$
15:  $w_{2}\leftarrow\partial_{s}\bar{\alpha}_{s}-\bar{\alpha}_{s}\,w_{1}$;  $w_{3}\leftarrow-\bar{\gamma}\,w_{1}$
16:  return $w_{1}\,\bar{x}_{s}+w_{2}\,\hat{z}+w_{3}\,x_{t}$
17:
@@ 4.1 GLASS Transitions
which defines the GLASS transition. This is a large family of transitions where $\rho$ controls the similarity between $X_{t}$ and $X_{t^{\prime}}$. In particular, it includes the most important example of DDPM transitions (see section A.2 for proof): {myframe}
@@ 4.2.1 GLASS denoiser
In this context, it is instructive to think of $x_{t}$ as a noisy measurement of a parameter $z$. The “standard” denoiser $D_{t}$ represents the mean of data $z$ given one Gaussian measurement $x_{t}$, while the GLASS denoiser $D_{\mu,\Sigma}$ represents the mean of data $z$ given two Gaussian measurements $(x_{t},x_{t^{\prime}})$ of $z$. Our core idea is that we can effectively “summarize” two Gaussian measurements $(x_{t},x_{t^{\prime}})$ into a single variable via the transformation
@@ 4.2.2 GLASS velocity field
This theorem shows that any flow matching or diffusion model contains an “inner” flow matching model $u_{s}(\bar{x}_{s}|x_{t},t)$ that allows to sample GLASS transitions. By proposition 2, no further training is required. As this result relies on the idea of using the sufficient statistic of Gaussian measurements to infer a latent $z$, we coin these flows Gaussian Latent Sufficient Statistic (GLASS) Flows.
4.3 Implementation
In algorithm 1, we describe pseudocode to sample a transition with GLASS Flows. We discuss here other aspects relevant for the implementation. To generate a data point $X_{1}\sim p_{\rm{data}}$, we set the number $K$ of transitions and transition times $0\leq t_{0}<t_{1}<\dots<t_{K}=1$ and initialize $X_{0}=X_{t_{0}}\sim\mathcal{N}(0,I_{d})$. For every $k=0,\cdots,K-1$, we sample the transition from $X_{t_{k}}$ to $X_{t_{k+1}}$ using algorithm 1. The number of function evaluations is $K\cdot M$ where $M$ is the number of simulation steps to simulate eq. 22. For $K=1$, only one transition, we recover standard flow matching as the conditioning for $t_{0}=0$ is simply ignored. Further, for $M=1$, one simulation step for the inner transition, we recover DDIM sampling (Song et al., 2020a) (see section B.2 for a derivation). Finally, all current large-scale flow matching models use classifier-free guidance (CFG) (Ho & Salimans, 2022) to condition on a prompt $c$. To use CFG with GLASS Flows, we treat the classifier-free guidance vector field $u_{t}^{w}(x|c)=(1+w)u_{t}(x|c)-wu_{t}(x)$ as the ground truth vector field for the same weight $w\geq 0$, i.e. all calculations are done with this vector field.
@@ 5 Related Work
We discuss the most closely related work in this section and refer to appendix C for an extended discussion of other related methods. GLASS Flows operate in discrete-time, leveraging an underlying continuous-time model. Discrete-time diffusion (Sohl-Dickstein et al., 2015) appeared prior to continuous-time diffusion, but such models are parameterized as 1st order approximations of the same ODE/SDE and are not qualitatively different. GLASS Flows instead consider sampling from latent Gaussian transitions for arbitrarily distant times. Transition Matching (Shaul et al., 2025) modifies pre-training and network architectures to sample transitions via flows, while our method focuses on inference-time modification post-training. The DTM supervision process in Transition Matching (see (Shaul et al., 2025, equ. (10))) corresponds to a GLASS transition with $\rho=1$.
Inference-time reward alignment methods are reviewed in (Uehara et al., 2025). They can be categorized into single particle (i.e. guidance) and more general multi-particle methods (i.e. Sequential Monte Carlo (SMC) and search). Guidance such as (Chung et al., 2022; Song et al., 2023b; Ye et al., 2024; Yu et al., 2023; Bansal et al., 2023; He et al., 2023; Song et al., 2023a; Feng et al., 2025) aims to approximate the difference between an existing velocity and an optimal velocity trained with a reward. In addition to guidance, source-based methods specific to flows keep the velocity fixed while altering the input noise distribution (Ben-Hamu et al., 2024; Eyring et al., 2024; Wang et al., 2025). Multi-particle methods evolve a population of particles such as SMC (Singhal et al., 2025; Skreta et al., 2025; Wu et al., 2023a; Mark et al., 2025) and search (Li et al., 2025; Zhang et al., 2025).

### v2

@@ 1 Introduction
Flow matching and diffusion models have revolutionized the generation of images, videos, and many other data types (Lipman et al., 2022; Albergo et al., 2023; Liu et al., 2022; Song et al., 2020b; Ho et al., 2020). They convert Gaussian noise into realistic images or videos by simulating an ordinary or stochastic differential equation (ODE/SDE). Trained on large web-scale datasets, these models can generate highly realistic images or videos at unprecedented quality. Due to diminishing returns of pre-training these models, many recent works propose methods to improve models at inference-time, i.e. by optimizing additional objectives commonly referred to as rewards (Uehara et al., 2025). These reward alignment algorithms allow to enhance text-to-image alignment (Zhang et al., 2025), solve inverse problems (Chung et al., 2022; He et al., 2023), and improve molecular design (Li et al., 2025b). However, as these algorithms achieve higher performance at the expense of more compute, efficiency remains a major challenge for deploying reward alignment algorithms.
@@
$p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t})=\mathbb{P}[X_{t^{\prime}}=x_{t^{\prime}}|X_{t}=x_{t}],\quad(x_{t},x_{t^{\prime}}\in\mathbb{R}^{d},0\leq t<t^{\prime}\leq 1)$ (1)
where $p_{t^{\prime}|t}$ is called the transition kernel. Many reward alignment algorithms rely on sampling from $p_{t^{\prime}|t}$. For example, search methods use samples $X_{t^{\prime}}\sim p_{t^{\prime}|t}(\cdot|x_{t})$ as branches of a search tree (for ODE sampling, there would be only one branch). This creates a dilemma: So far, it is not known how to obtain samples $X_{t^{\prime}}\sim p_{t^{\prime}|t}(\cdot|x_{t})$ using ODEs. Therefore, one has to switch from ODE to SDE sampling, losing efficiency in order to use an alignment algorithm that is meant to increase it.
@@
In this work, we present a method to sample from transitions $p_{t^{\prime}|t}$ using ODEs: Gaussian Latent Sufficient Statistic (GLASS) Flows. GLASS Flows combine (1) the high efficiency of ODEs with (2) the controllable stochastic evolution characteristic of SDEs. GLASS Flows construct an “inner flow matching model” to sample from $p_{t^{\prime}|t}$ (see fig. 1). Crucially, this inner flow matching can be easily obtained from pre-trained flow matching models without any fine-tuning - this transformation relies on the concept of a sufficient statistic, a fundamental tool in theoretical statistics (Fisher, 1922). Hence, GLASS Flows are a simple plug-in for any algorithm relying on SDE sampling. We apply GLASS Flows to reward alignment and improve the state-of-the-art performance in text-to-image generation. To summarize, we make the following contributions:
@@
    We show significant performance improvements for text-to-image generation at zero cost by plugging GLASS Flows into Sequential Monte Carlo and reward guidance procedures.
@@ 2 Background and Motivation
In this section, we introduce the necessary background on flow and diffusion models. We follow the flow matching (FM) framework (Lipman et al., 2022; Albergo et al., 2023; Liu et al., 2022), yet everything applies similarly to diffusion models (see section B.4). We denote data points with $z\in\mathbb{R}^{d}$ and the data distribution with $p_{\rm{data}}$. Here, $t=0$ corresponds to noise ($\mathcal{N}(0,I_{d})$) and $t=1$ to data $(p_{\rm{data}})$. To noise data $z\in\mathbb{R}^{d}$, we use a Gaussian conditional probability path $p_{t}(x_{t}|z)$:
@@
where $\alpha_{t},\sigma_{t}\geq 0$ are schedulers with $\alpha_{0}=\sigma_{1}=0$ and $\alpha_{1}=\sigma_{0}=1$ and $\alpha_{t}$ (resp. $\sigma_{t}$) strictly monotonically increasing (resp. decreasing) and continuously differentiable. With $z\sim p_{\rm{data}}$ random, this induces a marginal probability path $p_{t}(x_{t})=\mathbb{E}_{z\sim p_{\rm{data}}}[p_{t}(x_{t}|z)]$ which interpolates Gaussian noise $p_{0}=\mathcal{N}(0,I_{d})$ and data $p_{1}=p_{\rm{data}}$. FM models learn the marginal vector field:
@@ 3 Motivation: Efficient Transitions for Reward Alignment
Sequential Monte Carlo (SMC) methods (Wu et al., 2023a; Singhal et al., 2025; Skreta et al., 2025) use a transition kernel $p_{t^{\prime}|t}$ as a proposal distribution. They evolve $K$ particles $x_{t}^{k}$ via
@@
Search methods (Li et al., 2025b; Zhang et al., 2025) consider DDPM sampling as a rollout of a search tree with branches coming from samples from $p_{t^{\prime}|t}$. Beyond sampling branches of the search tree, search methods use approximations of the value function (Li et al., 2025b) defined via
@@
As sampling from $p_{1|t}(z|x_{t})$ is only possible with the SDE so far and therefore inefficient, most search methods use approximations of this function (Li et al., 2025b; Zhang et al., 2025). Similarly, approximations of the value function $V_{t}(x_{t})$ can also be used to define potentials in SMC procedures.
@@ 4 GLASS Flows
we get samples from the transition kernel $p_{t^{\prime}|t}$ at $s=1$ for an appropriate initial distribution $p_{\text{init}}$. Note that with this approach, we achieve stochasticity by sampling the inner initial condition $\bar{X}_{0}$, while the subsequent evolution follows a deterministic ODE. In contrast, SDE transitions have deterministic initial conditions but the increments are stochastic. We present a simple algorithm to obtain $u_{s}(\bar{x}_{s}|x_{t},t)$ that we explain in this section (see algorithm 1).
@@
2:  $u\leftarrow{\color[rgb]{0,0,0}u_{t}(x_{t})}$ $\triangleright$ neural net call
3:  return $\frac{1}{\dot{\alpha}_{t}\,\sigma_{t}-\alpha_{t}\,\dot{\sigma}_{t}}(\sigma_{t}u-\dot{\sigma}_{t}x_{t})$
@@
6:  If $s=0$: return $D_{t}(x_{t})$
7:  $S(\mathbf{x})\leftarrow\frac{\mu^{T}\Sigma^{-1}}{\mu^{T}\Sigma^{-1}\mu}[x_{t},\bar{x}_{s}]^{T}$
8:  $t^{*}\leftarrow g^{-1}((\mu^{T}\Sigma^{-1}\mu)^{-1})$
9:  return $D(\alpha_{t^{\star}}S(\mathbf{x}),t^{\star})$
10:
@@
12:  $\Sigma\leftarrow\begin{bmatrix}\sigma_{t}^{2}&\sigma_{t}^{2}\bar{\gamma}\\ \sigma_{t}^{2}\bar{\gamma}&\bar{\sigma}_{s}^{2}+\bar{\gamma}^{2}\sigma_{t}^{2}\end{bmatrix},\mu\leftarrow\begin{bmatrix}\alpha_{t}\\ \bar{\alpha}_{s}+\bar{\gamma}\alpha_{t}\end{bmatrix}$
@@
14:  $w_{1}\leftarrow\frac{\partial_{s}\bar{\sigma}_{s}}{\bar{\sigma}_{s}}$$w_{2}\leftarrow\partial_{s}\bar{\alpha}_{s}-\bar{\alpha}_{s}\,w_{1}$$w_{3}\leftarrow-\bar{\gamma}w_{1}$
15:  return $w_{1}\,\bar{x}_{s}+w_{2}\,\hat{z}+w_{3}\,x_{t}$
16:
@@ 4.1 GLASS Transitions
which defines the GLASS transition. This is a large family of transitions where $\rho$ controls the similarity between $X_{t}$ and $X_{t^{\prime}}$. It includes the important example of DDPM transitions: {myframe}
@@
See section A.2 for a proof. Note that $\rho$ defined like this is a valid correlation coefficient (i.e. $|\rho|\leq 1$) because $\frac{\sigma_{t^{\prime}}}{\sigma_{t}}<1$ and $\frac{\alpha_{t}}{\alpha_{t^{\prime}}}<1$ by monotonicity of the schedulers.
@@ 4.2.1 GLASS denoiser
Here, it is instructive to think of $x_{t}$ as a noisy measurement of a parameter $z$. The “standard” denoiser $D_{t}$ represents the mean of $z$ given one Gaussian measurement $x_{t}$, while the GLASS denoiser $D_{\mu,\Sigma}$ represents the mean of $z$ given two Gaussian measurements $(x_{t},x_{t^{\prime}})$. Our core idea is that we can effectively “summarize” two measurements $(x_{t},x_{t^{\prime}})$ into a single variable via the transformation
@@ 4.2.2 GLASS velocity field
This theorem shows that any flow matching or diffusion model contains an “inner” flow matching model $u_{s}(\bar{x}_{s}|x_{t},t)$ that allows to sample GLASS transitions. By proposition 2, no further training is required. As this result relies on the idea of using the sufficient statistic of Gaussian measurements to infer a latent $z$, we coin these flows Gaussian Latent Sufficient Statistic (GLASS) Flows. In algorithm 1, we describe pseudocode to sample a transition with GLASS Flows. Note that the reparameterizations and $2\times 2$ matrix inversions are negligible compared to neural network evaluations. Therefore, the complexity of algorithm 1 is governed by the number of function evaluations of the pre-trained velocity field $u_{t}(x)$, i.e. the number of simulation steps $M$.
Sampling with GLASS Flows. To generate a data point $X_{1}\sim p_{\rm{data}}$, we set the number $K$ of transitions and transition times $0\leq t_{0}<t_{1}<\dots<t_{K}=1$ and initialize $X_{0}=X_{t_{0}}\sim\mathcal{N}(0,I_{d})$. For every $k=0,\cdots,K-1$, we sample the transition from $X_{t_{k}}$ to $X_{t_{k+1}}$ using algorithm 1 (set $t=t_{k}$ and $t^{\prime}=t_{k+1}$) with a choice of a correlation parameter $\rho$ that we can choose freely (note that it can also vary across transitions). Assuming no discretization error and perfect training, if $X_{t_{k}}$ is distributed according to the probability path, also the next step will have marginals specified by the probability path by theorem 1:
$X_{t_{k}}\sim p_{t_{k}}\quad\Rightarrow\quad X_{t_{k+1}}\sim p_{t_{k+1}}$
In particular, it will hold that $X_{t_{K}}=X_{1}\sim p_{1}=p_{\rm{data}}$, i.e. the endpoint is a valid sample from the desired distribution. This preservation of marginals holds for any $\rho$ (not limited to the one corresponding to DDPM transitions). Therefore, GLASS is a novel sampling scheme resulting in Markov chains preserving the marginals of a pre-trained flow or diffusion models. The total number of function evaluations is $K\cdot M$. For $K=1$, only one transition, we recover standard flow matching as the conditioning for $t_{0}=0$ is simply ignored. Further, for $M=1$, one simulation step for the inner transition, we recover DDIM sampling (Song et al., 2020a) (see section B.2 for a derivation).
Implementation. For numerical stability, we need to account for the cases when $s=0$ in algorithm 1. We derive this edge case in section B.3. In section B.3, we also discuss other techniques to make the implementation numerically stable. Further, all current large-scale flow matching models use classifier-free guidance (CFG) (Ho & Salimans, 2022) to condition on a prompt $c$. To use CFG with GLASS Flows, we treat the classifier-free guidance vector field $u_{t}^{w}(x|c)=(1+w)u_{t}(x|c)-wu_{t}(x)$ as the ground truth vector field for the same weight $w\geq 0$, i.e. all calculations are done with this vector field. Finally, it is well-known that there are many equivalent parameterizations of the vector field $u_{t}$ (e.g. via the score function or directly via the denoiser) and also diffusion models in discrete time. We discuss in section B.4 how to construct GLASS Flows with these alternative parameterizations. Further, we provide a minimal implementation of algorithm 1 at github.com/PeterHolderrieth/glass_flows_tutorial.
@@ 5 Related Work
We discuss the most closely related work in this section and refer to appendix C for an extended discussion of other related methods. GLASS Flows operate in discrete-time, leveraging an underlying continuous-time model. Discrete-time diffusion (Sohl-Dickstein et al., 2015) appeared prior to continuous-time diffusion, but such models are parameterized as 1st order approximations of the same ODE/SDE and are not qualitatively different. GLASS Flows instead consider sampling from latent Gaussian transitions for arbitrarily distant times. Recently, discrete-time transitions in FM models were also studied in Transition Matching (Shaul et al., 2025). In fact, the DTM supervision process in Transition Matching (see (Shaul et al., 2025, equ. (10))) corresponds to a GLASS transition with $\rho=1$ (see appendix C for detailed discussion). However, note that Transition Matching (Shaul et al., 2025) modifies pre-training and network architectures via patch approximations to sample transitions via flows, while our method focuses on inference-time modification post-training. Therefore, GLASS Flows and TM address different problems and lead to different models that are theoretically related, yet practically different.
Inference-time reward alignment methods are reviewed in (Uehara et al., 2025). They can be categorized into single particle (i.e. guidance) and more general multi-particle methods (i.e. Sequential Monte Carlo (SMC) and search). Guidance such as (Chung et al., 2022; Song et al., 2023b; Ye et al., 2024; Yu et al., 2023; Bansal et al., 2023; He et al., 2023; Song et al., 2023a; Feng et al., 2025) aims to approximate the difference between an existing velocity and an optimal velocity trained with a reward. In addition to guidance, source-based methods specific to flows keep the velocity fixed while altering the input noise distribution (Ben-Hamu et al., 2024; Eyring et al., 2024; Wang et al., 2025). Multi-particle methods evolve a population of particles such as SMC (Singhal et al., 2025; Skreta et al., 2025; Wu et al., 2023a; Mark et al., 2025; He et al., 2025) and search (Li et al., 2025b; Zhang et al., 2025).
Recently, Chen et al. (2025b) propose Training-free Augmented Dynamics (TADA) introducing training-free improvements of diffusion models using a similar mathematical principle as in this work. Specifically, they show that several recently proposed diffusion models with augmented state spaces (Dockhorn et al., 2021; Chen et al., 2023) can be recovered from a pre-trained FM or diffusion model using a Gaussian conjugacy/sufficient statistic argument (Chen et al., 2025b, Proposition 3.1). Further, this principle can be extended to state spaces augmented with more than $2$ variables. This allows them to accelerate sampling significantly. While we design a different method designed to get fast stochastic transition samplers for reward alignment, both (Chen et al., 2025b) and this work use the same mathematical principles to derive their respective algorithms. We discuss this in more detail in appendix C.
Finally, reward fine-tuning methods based on GRPO (Xue et al., 2025; Li et al., 2025a; Liu et al., ), stochastic optimal control (Liu et al., ; Domingo-Enrich et al., 2024), DPO (Wallace et al., 2024) or other reinforcement learning approaches aim to achieve the same goal as this work, i.e. to align a diffusion model with a reward function $r$. GLASS has 2 different important synergies with these models: (1) Many of these algorithms require the use of DDPM/SDE sampling for exploration during training (Liu et al., 2025; Xue et al., 2025; Li et al., 2025a; Domingo-Enrich et al., 2024). However, this is very inefficient - as discussed in this work. Therefore, one could potentially accelerate reward fine-tuning methods via GLASS Flows by replacing the slow SDE sampling with GLASS Flows. (2) One can also apply inference-time scaling with GLASS Flows to fine-tuned models. Many of these models learn a FM model of reward-tilted distribution. Hence, GLASS Flows is equally valid to be applied to reward fine-tuned models and can be used to improve these models as well. Therefore, both approaches complement each other.